OpenStudy (chris215):

Find f[g(x)] if f(x) = x^4 + 1 and g(x) = x^2.

OpenStudy (chris215):

I got x^8+1

OpenStudy (solomonzelman):

Yes, correct:)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle f(\color{red}{x})=\color{red}{x}^4+1 }\) \(\color{#000000 }{ \displaystyle f[\color{red}{g(x)}]=\left(\color{red}{x^2}\right)^4+1 }\) \(\color{#000000 }{ \displaystyle f[g(x)]=x^{2\times 4}+1 =x^8+1 }\)